OpenStudy (arbyj):

x-2y=-3 in slope intercept form: y= -1/2x+2 can someone check my answer?

pooja195 (pooja195):

That' incorrect can you show me how you got that?

random231 (random231):

yeah your answer isnt right!

OpenStudy (arbyj):

i just realized i neglected the -3 :p

pooja195 (pooja195):

Ok x-2y=-3 Step 1 Subtract x from both sides -x-x-2y=-3-x What do we ahve now?

pooja195 (pooja195):

*have

OpenStudy (arbyj):

3

OpenStudy (arbyj):

wait nevermind

pooja195 (pooja195):

\[\huge~\rm~\bf −2y=−x−3 \]

OpenStudy (arbyj):

-4

pooja195 (pooja195):

to isolate y just divide everything by 2 what so we get? -2/-2=1 so just y y=_____x+_____

OpenStudy (arbyj):

-